OpenStudy (clamin):

ONE MEDAL!! Boyle's Law is stated by the formula: PV=pv Find V when P=220, p=25, and v=440. a.12.5 b.50 c.100 d.200

OpenStudy (aayushi.somani):

220*V=25*440 so V=50

OpenStudy (usukidoll):



OpenStudy (usukidoll):